You only have nickels and dimes. You have 5 times more dimes then nickels. You also have $1.35 more in dimes then you do nickles. How many dimes and nickels do you have?
Is this logical? It was a pop question in algebra. I have worked every solution I know.
Please and thank you,
James

Found 2 solutions by mananth, Alan3354:
Answer by mananth(16949)

nickels =x
dimes = 5x
10*5x-5*x=135
45x=135
/45
x=3 the number of nickels
dimes = 5*x= 5*3 = 15
CHECK
15*10-3*5= 135
=$1.35

Answer by Alan3354(69443)

You only have nickels and dimes. You have 5 times more dimes then nickels. You also have $1.35 more in dimes then you do nickles. How many dimes and nickels do you have?
Is this logical? It was a pop question in algebra. I have worked every solution I know.
---------------
5 times more is 6 times as many.
d = 6n
10d = 5n + 135
--------
10*6n = 5n + 135
60n = 5n + 135
55n = 135
No integer solution.
----------------------
5 times more is not 5 times as many.
100% more is 1 times more = 2 times as many, etc.
The problem is written in bad English.